# A Survey on Point-of-Interest Recommendation: Models, Architectures, and Security

### v1

@@ V-B Decentralized-based Architecture
TABLE III: Related works in different application sub-tasks. We provide detailed information on the names of journals/conferences, the specific generative techniques used, the particular application sub-tasks addressed, and the datasets used for evaluation.
@@
LCLCARS [64] KDD 2013 LDA Location-Content-Aware Recommender Foursquare, and Douban Event
LCARS [16] TOIS 2014 LDA Spatial Item Recommender Douban Event, and Foursquare
@@
TRM [65] ACM MM 2015 LDA POI Recommendation Foursquare, and Twitter
@@
SPORE [69] ICDE 2016 LDA Spatial Item Recommendation Foursquare, Twitter [168], and Synthetic Dataset
@@
TPM [74] TIST 2018 LDA Spatial Item Recommendation Foursquare, Twitter [168], and Synthetic Dataset
@@
LSMA [84] IJGI 2022 LSTM Next POI Recommendation Foursquare-Charlotte (CHA) [169], Foursquare-NYC
@@
CAFPR [98] Applied SC 2023 Transformer POI Recommendation Foursquare-TKY, CA, Budapest [170], Melbourne [170], Magic k [170]
@@
LLMmove [46] Arxiv 2024 LLM POI Recommendation NYC [171], TKY [171]
Li et al. [47] SIGIR 2024 LLM POI Recommendation NYC [171], TKY [171], Gowalla-CA
@@
DCPR [136] arXiv 2024 Diffusion POI Recommendation Foursquare, and Weeplace
@@
In this part, we illustrate the decentralized architecture for POI recommendation. While centralized architectures have excelled in accuracy, they require substantial computational resources during the training phase. To address these challenges, researchers have introduced decentralized architectures like [54, 27, 140]. To be specific, Long et al. [54] propose an innovative method for decentralized POI recommendations called Model-Agnostic Collaborative Learning for decentralized POI recommendation, which empowers users to tailor their own model architectures, such as adjusting dimensions and the quantity of hidden layers. To address the scarcity of user data on devices, MAC [54] introduces a strategy to pre-determine collaborators based on physical proximity, preferences at the category level, and social connections. For efficiently and securely integrating insights from these chosen collaborators, MAC implements a knowledge distillation approach with a focus on maximizing mutual information. In MAC, rather than exchanging sensitive models or gradients, clients selectively share their informed decisions concerning a predetermined reference dataset.
However, current collaborative learning systems commonly rely on a single reference dataset shared among all users, which may not cater to individual preferences, potentially hindering knowledge exchange and overall performance. Users may struggle to provide relevant insights on items outside their scope of interest, further complicating the collaborative process. To bridge this gap, Zheng et al. [140] propose the Adaptive Reference Data for Decentralized Collaborative Learning framework. DARD tailors adaptive reference datasets for enhanced user collaboration by creating a desensitized public reference pool using transformation and probability data generation techniques. For each user, DARD dynamically selects personalized reference data through loss tracking and influence function analysis. Individual models are trained using private data and collaboratively with spatial and contextual neighbors. During collaboration, users exchange soft decisions based on a blend of their personalized adaptive reference data, fostering more effective collaboration and personalized recommendations.
@@ V-C Federated Learning-based Architecture
Figure 5: Data poisoning (Left) and model poisoning (Right)
FL operates on the principle of training machine learning models without requiring access to individual users’ raw private data [142, 143, 144, 145, 146, 29, 147, 148, 149, 150]. An earlier pioneering work [142] introduces a cross-domain POI framework that incorporates federated learning techniques and privacy safeguards. By leveraging data from a related domain in user interest profiling, it aims to mitigate the cold-start issue. Additionally, the framework employs federated learning to locally analyze users’ historical data and encrypts latent feature distributions to ensure user privacy while transferring knowledge. The method aims to solve challenges like the sensitive and sparse nature of POI data and the cold-start dilemma stemming from data scarcity, which hampers recommendation effectiveness.
Another earlier work [148] introduces PREFER, a novel federated learning framework optimized for POI recommendation with edge acceleration. PREFER divides the recommendation process into distinct segments. Initially, for privacy preservation, users develop individualized recommendation models locally and exchange multi-dimensional user-agnostic parameters instead of sharing check-in data. Subsequently, to enhance recommendation speed, these decentralized parameters are consolidated on edge servers situated near users (like base stations), as opposed to distant cloud servers. Besides, Huang et al. [143] present a novel geographical Point of Interest recommendation approach based on federated learning. The POI recommendation task is framed as a matrix factorization optimization challenge, with the application of singular value decomposition for matrix decomposition. By demonstrating the non-convex nature of the optimization problem, the framework integrates stochastic gradient descent within SVD and devises a FL framework for concurrently addressing the POI recommendation task. Within our FL strategy, solely the computed gradient details are transmitted from users to the FL server, while users independently handle their rating and geographic preference data on their personal devices to ensure privacy during communication.
FedPOIRec [29] is a privacy-focused framework that ensures local data remains on the user’s device while updates are aggregated blindly by a parameter server. To enhance personalization, FedPOIRec enables users to exchange learned parameters with their friends, facilitating knowledge transfer. To further protect privacy, we introduce a protocol that integrates the preferences of a user’s friends after federated computation. This protocol leverages the properties of the CKKS fully homomorphic encryption scheme to ensure privacy during this integration process.
@@ VI Security
VI-A Data Integrity Threats
@@ VI-B User Privacy Protection
To address privacy concerns, various cryptographic techniques and privacy-preserving mechanisms are essential for safeguarding sensitive information in POI recommendation systems. Decentralized secure computation [60, 61] enables multiple parties to collaboratively compute a function over their inputs while keeping them private. Chen et al. [60] propose a decentralized matrix factorization framework for POI recommendation, addressing privacy and computational issues by enabling decentralized training on user devices. This approach preserves users’ rating data and enhances recommendation accuracy. Furthermore, Chen et al. [61] introduce the PriRec framework, which maintains users’ private data on their devices while employing local differential privacy techniques and secure decentralized gradient descent to protect model privacy, achieving comparable recommendation accuracy.
@@
Additionally, homomorphic encryption [63, 29] is also commonly used for privacy-preserving POI recommendation. It allows computations on encrypted data without decryption, ensuring sensitive information remains secure. Figure 7 illustrates the workflow of homomorphic encryption, which conceals the underlying plaintext during processing. Li et al. [63] develop a privacy-preserving framework for POI recommendation by utilizing homomorphic encryption. They address the cold start problem prevalent in POI recommendation systems, which arises due to the sparsity of user-location check-in data. Their framework leverages partially homomorphic encryption to enable trust-oriented recommendations while ensuring that sensitive user check-in and trust data remain private. By employing offline encryption and parallel computing, Li et al. [63] ensure that their protocols efficiently protect the private data of all parties involved in the recommendation process. They also prove that their proposed protocols are secure against semi-honest adversaries, and their experiments on both synthetic and real datasets demonstrate the feasibility of privacy-preserving recommendations with acceptable computation and communication costs.
@@ VII Future Directions
As POI recommendation systems continue to evolve, vast opportunities for innovation will keep emerging across the three pivotal dimensions: models, architectures, and security. By integrating cutting-edge model structures, adopting flexible architectures, and reinforcing security measures, the future of POI recommendation systems are expected to be more adaptive, personalized, and resilient to emerging challenges. Below, we outline the key future directions that will shape the development of these systems across these essential aspects.
@@
Future POI recommendation models will move beyond traditional and deep learning techniques, incorporating various data modalities and more advanced methodologies to further improve accuracy and personalization. (1) LLM Agent-Powered POI Recommendation. The integration of Large Language Model (LLM) agents [172] into POI recommendation systems offers another approach to enhance personalization POI recommendation. LLM agents can interpret user preferences through natural language interactions, making real-time POI recommendations based on conversational queries and feedback [173, 174]. These agents can provide personalized POI suggestions by engaging users in dialogue to clarify needs and generate tailored recommendations. By incorporating contextual factors such as location, time, and user-specific constraints, LLM-powered POI recommendation systems offer dynamic, context-aware suggestions. (2) LLM-Powered Explainable POI Recommendation. Creating an explainable POI recommendation system via large language models is a promising avenue for leveraging advanced language processing capabilities to provide transparent and understandable POI recommendations. By utilizing these models, users can receive recommendations accompanied by explanations that clarify why a particular POI is suggested. This transparency enhances user trust and engagement by shedding light on the decision-making process behind each recommendation, ultimately improving the overall user experience of POI recommendation.
Future POI recommendation architecture is to ensure that decentralized systems can deliver the same level of personalization and accuracy as their centralized counterparts, all while maintaining the privacy benefits that decentralized architectures provide. Small Language Model-Powered Decentralized POI Recommendation. While the mainstream focus is on developing large language models, small language models are also gaining attention for their efficiency and ability to run on-device, offering privacy-preserving and real-time capabilities. The newly released Llama 3.2’s lightweight, language models (1B and 3B)  [^1] https://ai.meta.com/blog/llama-3-2-connect-2024-vision-edge-mobile-devices/ can revolutionize edge AI and significantly improve the accuracy of decentralized POI recommendation systems by processing real-time, contextual data directly on the device. These models can learn from user interactions, preferences, and location data locally by using the long contexts (up to 128K tokens), allowing for highly personalized recommendations without needing centralized data aggregation. By running on-device, the models can continually adapt to user behavior, leading to more precise recommendations based on immediate context, while preserving privacy. The emergence of new models and architectures has brought up new opportunities, consequently expanding the landscape of potential threats. Future POI recommendation systems need to handle new security issues. (1) Addressing Vulnerabilities in Emerging POI Recommender Systems. As novel techniques such as multi-modal data infusion, self-supervised learning, and LLM-based recommendation systems gain traction, they introduce new layers of complexity and new vulnerabilities. The security landscape of POI recommendation
systems will need to focus on understanding these emerging attack surfaces, as they differ significantly from traditional techniques. Future research should prioritize identifying and mitigating vulnerabilities unique to these advanced techniques, such as adversarial manipulation in multi-modal inputs or exploiting context-driven weaknesses in LLM-based systems. By developing robust security frameworks tailored to these innovative approaches, future POI recommendation systems can safeguard against both known and unforeseen threats, ensuring trust and reliability as these technologies evolve. (2) Privacy-Enhancing POI Recommendation based on Cryptography Techniques. Developing privacy-enhancing techniques for POI recommendation systems is crucial for protecting user data while maintaining recommendation quality. This involves implementing methods such as secure multiparty computation and zero-knowledge proof, ensuring user interactions and preferences remain confidential. By safeguarding sensitive information, these techniques can enhance user trust and encourage broader adoption of recommendation systems, thereby balancing privacy protection with system performance and ensuring that privacy-enhanced models do not compromise the accuracy and relevance of recommendations. By embracing these diverse research directions and leveraging cutting-edge technologies and methodologies, the field of POI recommendation is poised to revolutionize the way users discover and engage with points of interest in the digital landscape.

### v2

@@ V-B Decentralized Architecture
In this part, we illustrate the decentralized architecture for POI recommendation. While centralized architectures excel in accuracy, they require substantial computational resources during training. To address these challenges, researchers have introduced decentralized architectures like [124, 28, 125]. Specifically, Long et al. [124] propose Model-Agnostic Collaborative Learning (MAC), which allows users to tailor model architectures, such as adjusting dimensions and the number of hidden layers. To mitigate the scarcity of user data on devices, MAC [124] pre-determines collaborators based on physical proximity, preferences, and social connections. It integrates insights from collaborators using a knowledge distillation approach, maximizing mutual information, and selectively sh-
TABLE IV: Related works in various application sub-tasks. We provide details on journal/conference names, generative techniques used, specific sub-tasks addressed, and evaluation datasets.
@@
LCLCARS [40] KDD 2013 LDA Location-Content-Aware Recommendation Foursquare, and Douban Event
LCARS [17] TOIS 2014 LDA Spatial Item Recommendation Douban Event, and Foursquare
@@
TRM [41] CIKM 2015 LDA POI Recommendation Foursquare, and Twitter
@@
SPORE [45] ICDE 2016 LDA Spatial Item Recommendation Foursquare, Twitter, and Synthetic Dataset
@@
TPM [50] TIST 2018 LDA Spatial Item Recommendation Foursquare, Twitter, and Synthetic Dataset
@@
LSMA [60] IJGI 2022 LSTM Next POI Recommendation Foursquare-Charlotte (CHA), Foursquare-NYC
@@
CAFPR [75] Applied SC 2023 Transformer POI Recommendation Foursquare-TKY, CA, Budapest, Melbourne, Magic k
@@
SCL [169] Expert Systems 2025 SSL POI Recommendation Florence, Rome, Pisa, Edinburgh, and Toront
@@
LLMmove [112] CAI 2024 LLM POI Recommendation NYC, TKY
Li et al. [113] SIGIR 2024 LLM POI Recommendation NYC, TKY, Gowalla-CA
@@
DCPR [117] KDD 2024 Diffusion POI Recommendation Foursquare, and Weeplace
@@
aring informed decisions on a reference dataset rather than sensitive models or gradients. However, current collaborative learning systems rely on a single shared reference dataset, which may not align with individual preferences and hinder knowledge exchange. To address this, Zheng et al. [125] propose the Adaptive Reference Data for Decentralized Collaborative Learning (DARD), which creates adaptive reference datasets for improved collaboration. DARD uses transformation and probability data generation techniques to create a public reference pool, from which personalized reference data is selected dynamically for each user based on loss tracking and influence function analysis. Users train models with private data and collaborate with spatial and contextual neighbors, exchanging soft decisions based on their personalized reference data.
While both MAC and DARD have merits, they suit different scenarios. MAC is effective when a reference dataset is available, such as in cases of public check-in sequences shared under authorization. DARD is preferred when no reference data exists, as it creates adaptive datasets for individual preferences. Each method excels under specific conditions, and neither method universally outperforms the other.
@@ V-C Federated Learning-based Architecture
FL operates on the principle of training machine learning models without requiring access to individual users’ raw private data [127, 128, 129, 130, 131, 30, 132, 133, 134, 136]. An earlier work [127] introduces a cross-domain POI framework that incorporates federated learning and privacy safeguards. It leverages data from related domains to address the cold-start issue and employs federated learning to analyze users’ historical data locally, encrypting feature distributions to ensure privacy while transferring knowledge. This method addresses challenges like sensitive POI data and cold-start problems caused by data scarcity. Another work presents PREFER [133], a federated learning framework for POI recommendation with edge acceleration. Users develop individualized models locally and exchange multi-dimensional, user-agnostic parameters for privacy. These parameters are then consolidated on edge servers, enhancing recommendation speed. Huang et al. [128] propose a federated learning-based geographical POI recommendation approach, framing the task as a matrix factorization problem solved through singular value decomposition and stochastic gradient descent, with only gradients transmitted for privacy protection. In addition, a privacy-focused framework, named FedPOIRec [30], was proposed to integrate the preferences of a user’s friends after the federated computation. In this framework, data remains on the user’s device, and updates are aggregated by a parameter server, and it allows users to exchange learned parameters with friends to improve personalization by using a CKKS fully homomorphic encryption scheme to ensure privacy during this process.
Each method has its strengths depending on the scenario. The cross-domain POI framework [127] is useful when related domain data is available to address cold-start issues. PREFER [133] excels in environments requiring privacy and fast, localized recommendations using edge servers. Huang et al.’s [128] method is ideal when geographical data is crucial, while FedPOIRec [30] is best for applications relying on social connections for personalized recommendations. Each approach offers advantages under specific conditions.
@@ VI Security
VI-A System Vulnerability Analysis
@@
Figure 6: Data poisoning (Left) and model poisoning (Right)
@@ VI-B User Privacy Protection
To address privacy concerns, various cryptographic techniques and privacy-preserving mechanisms are essential for safeguarding sensitive information in POI recommendation systems. Decentralized secure computation [135, 143] enables multiple parties to collaboratively compute a function over their inputs while keeping them private. Chen et al. [135] proposed a decentralized matrix factorization framework for POI recommendation, addressing privacy and computational issues by enabling decentralized training on user devices. This approach preserves users’ rating data and enhances recommendation accuracy. Furthermore, Chen et al. [143] introduced the PriRec framework, which maintains users’ private data on their devices while employing local differential privacy techniques and secure decentralized gradient descent to protect model privacy, achieving comparable recommendation accuracy.
Additionally, homomorphic encryption [144, 30] is also commonly used for privacy-preserving POI recommendation. It allows computations on encrypted data without decryption, ensuring sensitive information remains secure. Figure 7 illustrates the workflow of homomorphic encryption, which conceals the underlying plaintext during processing. Li et al. [144] developed a privacy-preserving framework for POI recommendation by utilizing homomorphic encryption. They address the cold start problem prevalent in POI recommendation systems, which arises due to the sparsity of user-location check-in data. Their framework leverages partially homomorphic encryption to enable trust-oriented recommendations while ensuring that sensitive user check-in and trust data remain private. By employing offline encryption and parallel computing, Li et al. [144] ensured that their protocols efficiently protect the private data of all parties involved in the recommendation process. They also prove that their proposed protocols are secure against semi-honest adversaries, and their experiments on both synthetic and real datasets demonstrate the feasibility of privacy-preserving recommendations with acceptable computation and communication costs.
@@ VII Future Directions
As POI recommendation systems continue to evolve, vast opportunities for innovation will keep emerging across the three pivotal dimensions: models, architectures, and security. By integrating cutting-edge model structures, adopting flexible architectures, and reinforcing security measures, the future of POI recommendation systems is expected to be more adaptive, personalized, and resilient to emerging challenges. Below, we outline the key future directions that will shape the development of these systems across these essential aspects.
VII-A Models
@@
Future POI recommendation models will move beyond traditional and deep learning techniques, incorporating various data modalities and more advanced methodologies to further improve accuracy and personalization. (1) LLM Agent-Powered POI Recommendation. The integration of Large Language Model (LLM) agents [170] into POI recommendation systems offers another approach to enhance personalization POI recommendation. LLM agents can interpret user preferences through natural language interactions, making real-time POI recommendations based on conversational queries and feedback [171, 172]. These agents can provide personalized POI suggestions by engaging users in dialogue to clarify needs and generate tailored recommendations. By incorporating contextual factors such as location, time, and user-specific constraints, LLM-powered POI recommendation systems offer dynamic, context-aware suggestions. (2) LLMs-Powered Explainable POI Recommendation. Creating an explainable POI recommendation system via large language models is a promising avenue for leveraging advanced language processing capabilities to provide transparent and understandable POI recommendations. By utilizing these models, users can receive recommendations accompanied by explanations that clarify why a particular POI is suggested. This transparency enhances user trust and engagement by shedding light on the decision-making process behind each recommendation, ultimately improving the overall user experience of POI recommendation. (3) Semantic Alignment Between POI Models and Real Applications. Bridging semantic gaps between spatial data and real-world contexts enables POI models to understand POI spatial information. Integrating geographic hierarchies, proximity patterns, and spatial correlations among POIs (e.g., analyzing a park’s relation to transit hubs and neighborhoods) will improve POI recommendation contextual relevance, thereby improving POI recommendation accuracy.
VII-B Architectures
Future POI recommendation architectures will primarily be distributed, incorporating technologies such as decentralized and federated recommendation. In actual deployment, these approaches face efficiency bottlenecks, making it difficult to effectively meet the demands of large-scale systems. The following outlines several future research directions focusing on scalability, latency, and communication efficiency. (1) Scalability Enhancement in Distributed POI Recommendation. Improving the scalability of POI recommendation architectures is crucial for their effective deployment in large-scale systems. Current distributed systems struggle to accommodate an increasing number of computational nodes due to the overhead of large-scale data processing and inter-node coordination. The primary challenge lies in the need for efficient model training and result aggregation while maintaining computational efficiency. Future research could explore highly collaborative recommendation frameworks that dynamically balance computational workloads, optimize resource allocation, and simplify coordination among nodes to improve scalability and ensure efficient system expansion. (2) Latency Optimization for Real-Time POI Recommendation. Reducing latency is a fundamental requirement for achieving real-time POI recommendations, particularly in decentralized collaborative systems. These architectures rely on multiple distributed nodes to perform recommendation tasks, but frequent data synchronization, inter-node communication, and model parameter exchanges introduce significant delays, impacting responsiveness. To mitigate these issues, future research could focus on developing asynchronous communication protocols, parallel computation strategies, and lightweight update mechanisms to enhance real-time performance in decentralized POI recommendation systems. (3) Communication-Efficient Federated POI Recommendation. Federated recommendation models face significant challenges in handling the transmission of large-scale model parameters across distributed nodes [129, 131]. As model sizes increase, resource-constrained nodes may struggle to efficiently participate in global model training, limiting the scalability of federated learning. To address this challenge, future research could focus on designing communication-efficient aggregation algorithms, adaptive parameter compression techniques, and incentive-driven task distribution mechanisms to optimize the efficiency of federated POI recommendation and enable seamless collaboration across diverse computational environments.
VII-C Security
Future research needs to address the challenge of balancing security and efficiency in existing privacy-preserving POI recommendation systems, as traditional techniques such as federated learning, differential privacy, and cryptographic methods often encounter efficiency issues due to their high computational or communication costs. The following research directions are worth exploring. (1) Application-Oriented Privacy-Preserving POI Recommendation. Optimizing privacy-preserving POI recommendation systems for specific application scenarios is crucial for improving efficiency. Different privacy protection methods exhibit varying trade-offs between security and computational cost. Federated learning excels in distributed training environments by safeguarding user privacy while minimizing data transmission overhead. Cryptographic techniques, such as homomorphic encryption and secure multi-party computation, are better suited for small-scale inference tasks that require strong privacy guarantees but lower computational complexity. Additionally, emerging machine unlearning techniques [173] offer efficient solutions for large-scale training and inference tasks, making them promising candidates for privacy-preserving POI recommendation. Future research should focus on designing adaptive frameworks that dynamically select the most suitable privacy-preserving strategy based on task characteristics such as training mode, inference complexity, and data scale. (2) TEE-Based Privacy-Preserving POI Recommendation. Leveraging Trusted Execution Environments (TEEs) presents a promising direction for enhancing security while maintaining computational efficiency in POI recommendation. TEEs, such as Intel SGX [174] and Arm TrustZone [175], provide secure enclaves for executing privacy-sensitive computations with minimal performance overhead. However, they remain susceptible to side-channel attacks and hardware vulnerabilities. Future research should explore techniques to enhance TEE security, such as mitigating known attack vectors and integrating lightweight cryptographic enhancements. By improving the robustness of TEEs, privacy-preserving POI recommendation systems can achieve secure and efficient computation without significant trade-offs in performance. Additionally, emerging cross-domain technologies, such as data watermarking, could be integrated to further enhance both security and efficiency in privacy-preserving POI recommendation systems. By combining technological customization with hardware optimization, future systems can improve privacy protection effectiveness while maintaining efficiency.